# Shuffle Mamba: State Space Models with Random Shuffle for Multi-Modal Image Fusion

### v1

@@ Introduction
Random Channel Interactive Mamba Block.
In the RCIM block, we refer to the approach from (He et al. 2024a) to achieve lightweight feature interaction between different modalities. We use the split operation to divide modality features $\mathbf{F}_{k}^{m}$ and $\mathbf{F}_{k}^{p}$ into two halves based on the channel dimension, followed by complementary splicing. The exchanged features are then sent to their respective RM Blocks for processing. By repeating these steps, the global modality-specific features are initially fused.
Random Modal Interactive Mamba block.
@@
Testing with Monte Carlo averaging
Table 1: Quantitative comparison of pan-sharpening task on three datasets. Bold and underline show the best and second-best values, respectively. $\uparrow$ indicates that the larger the value, the better the performance, and $\downarrow$ indicates that the smaller the value, the better the performance.
@@
INNformer AAAI’22 41.6903 0.9704 0.0227 0.9514 47.3528 0.9893 0.0102 0.5479 30.5365 0.9225 0.0747 3.1142
SFINet ECCV’22 41.7244 0.9725 0.0220 0.9506 47.4712 0.9901 0.0102 0.5479 30.5901 0.9236 0.0741 3.0798
MSDDN TGRS’23 41.8435 0.9711 0.0222 0.9478 47.4101 0.9895 0.0101 0.5414 30.8645 0.9258 0.0757 2.9581
PanFlowNet ICCV’23 41.8548 0.9712 0.0224 0.9335 47.2533 0.9884 0.0103 0.5512 30.4873 0.9221 0.0751 2.9531
FAME AAAI’24 42.0262 0.9723 0.0215 0.9172 47.6721 0.9898 0.0098 0.5542 30.9903 0.9287 0.0697 2.9531
DISPNet AAAI’24 41.8768 0.9702 0.0221 0.9157 47.4529 0.9898 0.0111 0.5532 30.0426 0.9153 0.0776 3.2620
Pan-mamba arxiv’24 42.2354 0.9729 0.0212 0.8975 47.6453 0.9894 0.0103 0.5286 31.1740 0.9302 0.0698 2.8910
Ours - 42.3428 0.9734 0.0208 0.8840 47.9180 0.9903 0.0097 0.5109 31.4005 0.9327 0.0676 2.8098
Table 2: Evaluation of our method on real-world full-resolution scenes from the WV2 dataset. Bold and underline show the best and second-best values, respectively.
Metric SFIM Brovey GFPCA INNformer SFINet PanFlowNet FAME DISPNet Pan-mamba Ours
$D_{\lambda}$downarrow 0.1403 0.1026 0.1139 0.0995 0.1034 0.0966 0.0951 0.0944 0.0966 0.0941
$D_{S}$downarrow 0.1320 0.1409 0.1535 0.1305 0.1305 0.1274 0.1263 0.1264 0.1272 0.1266
QNR $\uparrow$ 0.7826 0.7728 0.7532 0.7858 0.7827 0.7910 0.7933 0.7938 0.7911 0.7939
@@
We incorporate stochastic factors into the random shuffle operation, necessitating the marginalization of these factors in the generation of the final fusion. However, the random shuffle method presents a theoretical challenge due to the exponentially large number of potential models, making precise averaging of their predictions infeasible. Drawing inspiration from dropouts (Srivastava et al. 2014), we approximate the expected value of the entire model through layered expectations. Therefore, the computation of the random shuffle during testing can be expressed as follows, where $\mathbf{RM}=RandomMamba$:
$\mathbf{R}{\mathbf{M}^{test}}(x)=\underset{\mathbf{\mathcal{S}}}{\mathop{\mathbb{missing}}{E}}\,[\mathbf{RM}(x,\mathbf{\mathcal{S}})]$ (13)
In fact, estimating $\mathbf{R}{{\mathbf{M}}^{test}}$ based on the aforementioned equation requires enumerating all possible shuffle results, which imposes a significant computational burden. Therefore, we employ Monte Carlo averaging to estimate its expectation:
Specifically, the input image is independently shuffled ${M}$ times, and then the ${M}$ outputs of $\mathbf{RM}$ are calculated. The mean of these outputs is computed to obtain the final estimate. When ${M}\to\infty$ is present, the Monte Carlo estimator closely approximates the true mean. Figure 4 illustrates the testing process we designed, which significantly reduces the actual testing time by incorporating multiple identical inputs in a mini-batch and utilizing GPUs for parallel computation.
Table 3: Quantitative comparison of MIF task on three datasets. bold and underline show the best and second-best values.
@@
PSLPT 0.888 0.548 0.373 0.815 0.675 0.502 0.432 0.810 0.850 0.359 0.325 0.933
EMFusion 0.943 0.685 0.783 1.221 1.190 0.552 0.475 1.266 0.885 0.665 0.692 1.212
MSRPAN 1.017 0.581 0.799 1.182 1.319 0.436 0.455 1.261 0.960 0.525 0.560 1.153
SwinFusion 1.642 0.703 0.683 0.725 1.537 0.522 0.545 0.579 1.678 0.744 0.720 0.684
Zero 0.950 0.635 0.774 1.162 0.768 0.320 0.582 1.199 1.046 0.582 0.681 1.180
U2Fusion 0.947 0.460 0.292 0.494 0.309 0.074 0.489 0.042 0.865 0.419 0.696 0.479
CDDFuse 1.481 0.650 0.765 1.227 1.589 0.526 0.530 1.224 0.995 0.786 0.719 1.169
Ours 1.491 0.797 0.741 1.256 1.580 0.546 0.592 1.330 1.470 0.820 0.747 1.240
@@
Table 4: Ablation for our methods on three datasets. The PSNR/SSIM/SAM/ERGAS values on benchmarks are reported. The best results are shown in bold.
@@
Core Operation RM w/o Random Shuffle 42.1460/0.9723/0.0211/0.9042 47.4284/0.9890/0.0102/0.5355 31.3915/0.9329/0.0676/2.8113
RCIM w/o Random Shuffle 42.2443/0.9729/0.0210/0.8919 47.6847/0.9899/0.0099/0.5292 31.3476/0.9325/0.0680/2.8230
RMIM w/o Random Shuffle 42.2136/0.9727/0.0211/0.8978 47.5787/0.9895/0.0100/0.5324 31.0971/0.9295/0.0704/2.9097
Loss Function
In accordance with established norms within this area, the loss function employed in our model for pan-sharpening is the L1 loss. In the MIF task, we use two input images and the fused image to calculate a composite loss function that includes L1 loss, SSIM loss, and gradient loss.
Experiments
Datasets and Benchmark
For the pan-sharpening task, we utilized datasets from WorldView-II (WV2), GaoFen2 (GF2), and WorldView-III (WV3), encompassing a variety of urban and natural scenes. We generated training samples following the Wald (Wald, Ranchin, and Mangolini 1997) protocol in the inaccessible ground truth. We conducted a thorough comparison of our proposed method against classical approaches and deep learning-based methods, including INNformer (Zhou et al. 2022a), SFINet (Zhou et al. 2022b), MSDDN (He et al. 2023), PanFlowNet (Yang et al. 2023), FAME (He et al. 2024b), DISPNet (Wang et al. 2024) and Pan-mamba (He et al. 2024a).
In the MIF task, we employed medical images from the Harvard Medical website, comprising pairs such as MRI-CT, MRI-PET, and MRI-SPECT. For this task, we compared our method with various deep learning-based multi-modal fusion techniques, including PSLPT (Wang, Deng, and Vivone 2024), EMFusion (Xu and Ma 2021), MSRPAN (Fu et al. 2021), SwinFusion (Ma et al. 2022), Zero (Lahoud and Süsstrunk 2019), U2Fusion (Xu et al. 2020), and CDDFuse (Zhao et al. 2023).
Implementation Details
All experiments were conducted using the PyTorch framework on two NVIDIA RTX 3060 GPUs. We trained for 500 epochs with a batch size of 4 for the pan-sharpening task and 200 epochs with a batch size of 1 for the medical image fusion (MIF) task. The network parameters were optimized using the Adam optimizer. The initial learning rate was set to $45e-$ 4, which was subsequently reduced to $85e-$ 8 using the CosineAnnealingLR scheduler over the specified epochs. For the pan-sharpening task, we randomly cropped training set images to obtain LRMS patches of size 32x32 and PAN images of size 128x128. For the MIF task, the training set images were cropped to 256x256.
Comparison with SOTA Methods
Pan-sharpening.
@@
The experimental results on three datasets are presented in Table 1. Reference metrics, including PSNR, SSIM, SAM, and ERGAS (Alparone et al. 2007), are used to evaluate the fusion effect. The results demonstrate that the proposed method outperforms the SOTA methods across all metrics. Notably, in PSNR metric, our method achieves improvements of 0.1047, 0.2727, and 0.2301 dB over Pan-mamba, which has the second-best performance. Better PSNR and SSIM evidence that the fusion results transfers more information from the original image and experiences less distortion. In the qualitative comparison, Figure 5 shows a representative sample from the WV3 dataset. Our method performs better on the MSE graph, indicating that the fusion result is closer to the ground truth. Compared to other methods, our approach achieves a more accurate restoration of spectral and spatial details, highlighting the advantages of our fusion technique.
@@
To further verify the generalization ability of our method in full-resolution scenes, we use three non-reference metrics, including ${{D}_{s}}$, ${{D}_{\lambda}}$ and QNR, to evaluate our method on the full WorldView-II dataset. Table 2 presents the experimental results. Our method outperforms the comparison methods across all results, demonstrating the strong adaptability of Shuffle Mamba in image fusion.
Medical Image Fusion.

### v2

@@ III-C2 Random Mamba Block
III-C3 Random Channel Interactive Mamba Block
In the RCIM block, we refer to the approach from [12] to achieve lightweight feature interaction between different modalities. We use the split operation to divide modality features $\mathbf{F}_{k}^{m}$ and $\mathbf{F}_{k}^{p}$ into two halves based on the channel dimension, followed by complementary splicing. The exchanged features are then sent to their respective RM Blocks for processing. By repeating these steps, the global modality-specific features are initially fused.
III-C4 Random Modal Interactive Mamba Block
@@
III-D Testing with Monte Carlo averaging
We incorporate stochastic factors into the random shuffle operation, necessitating the marginalization of these factors in the generation of the final fusion. However, the random shuffle method presents a theoretical challenge due to the exponentially large number of potential models, making precise averaging of their predictions infeasible. Drawing inspiration from dropouts [48], we approximate the expected value of the entire model through layered expectations. Therefore, the computation of the random shuffle during testing can be expressed as follows, where $\mathbf{RM}=RandomMamba$:
$\mathbf{R}{\mathbf{M}^{test}}(x)=\underset{\mathbf{\mathcal{S}}}{\mathop{\mathbb{E}}}\,[\mathbf{RM}(x,\mathbf{\mathcal{S}})]$ (13)
Estimating $\mathbf{R}{{\mathbf{M}}^{test}}$ based on the aforementioned equation requires enumerating all possible shuffle results, which imposes a significant computational burden. Therefore, we employ Monte Carlo averaging to estimate its expectation:
Specifically, the input image is independently shuffled M{M} times, and then the M{M} outputs of $\mathbf{RM}$ are calculated. The mean of these outputs is computed to obtain the final estimate. When ${M}\to\infty$ is present, the Monte Carlo estimator closely approximates the true mean. Figure 4 illustrates the test process we designed, which significantly reduces the actual testing time by incorporating multiple identical inputs in a mini-batch and utilizing GPUs for parallel computation.
TABLE I: Quantitative comparison of pan-sharpening task on three datasets. The best results are highlighted in bold. The second-best results are highlighted in underline. $\uparrow$ indicates that the larger the value, the better the performance, and $\downarrow$ indicates that the smaller the value, the better the performance.
@@ III-D Testing with Monte Carlo averaging
SFIM 34.1297 0.8975 0.0439 2.3449 36.9060 0.8882 0.0318 1.7398 21.8212 0.5457 0.1208 8.9730
Brovey 35.8646 0.9216 0.0403 1.8238 37.2260 0.9034 0.0309 1.6736 22.5060 0.5466 0.1159 8.2331
IHS 35.6376 0.9176 0.0423 1.8774 37.7974 0.9026 0.0218 1.3720 22.5608 0.5470 0.1217 8.2433
GS 35.2962 0.9027 0.0461 2.0278 38.1754 0.9100 0.0243 1.5336 22.5579 0.5354 0.1266 8.3616
GFPCA 34.5581 0.9038 0.0488 2.1411 37.9443 0.9204 0.0314 1.5604 22.3344 0.4826 0.1294 8.3964
SFNLR 33.8636 0.8957 0.0481 2.2415 36.2367 0.8850 0.0363 1.8611 21.7647 0.5636 0.1494 9.0962
LRTCFPan 34.7756 0.9112 0.0426 2.0075 36.9253 0.8946 0.0332 1.7060 22.1574 0.5735 0.1380 8.6796
MSDCNN 41.3355 0.9664 0.0242 0.9940 45.6847 0.9827 0.0135 0.6389 30.3038 0.9184 0.0782 3.1884
INNformer 41.6903 0.9704 0.0227 0.9514 47.3528 0.9893 0.0102 0.5479 30.5365 0.9225 0.0747 3.1142
SFINet 41.7244 0.9725 0.0220 0.9506 47.4712 0.9901 0.0102 0.5479 30.5901 0.9236 0.0741 3.0798
MSDDN 41.8435 0.9711 0.0222 0.9478 47.4101 0.9895 0.0101 0.5414 30.8645 0.9258 0.0757 2.9581
PanFlowNet 41.8548 0.9712 0.0224 0.9335 47.2533 0.9884 0.0103 0.5512 30.4873 0.9221 0.0751 2.9531
FAME 42.0262 0.9723 0.0215 0.9172 47.6721 0.9898 0.0098 0.5542 30.9903 0.9287 0.0697 2.9531
DISPNet 41.8768 0.9702 0.0221 0.9157 47.4529 0.9898 0.0111 0.5532 30.0426 0.9153 0.0776 3.2620
Pan-mamba 42.2354 0.9729 0.0212 0.8975 47.6453 0.9894 0.0103 0.5286 31.1740 0.9302 0.0698 2.8910
ARConv 41.8281 0.9704 0.0222 0.9395 47.3699 0.9887 0.0102 0.5492 30.8156 0.9251 0.0717 3.0114
Ours 42.3428 0.9734 0.0208 0.8840 47.9180 0.9903 0.0097 0.5109 31.4005 0.9327 0.0676 2.8098
III-E Loss Function
In accordance with established norms within this area, the loss function used in our model for pan-sharpening is the L1 loss. Assume the two input images of the model be represented as $\mathbf{M}_{in}$} and $\mathbf{P}_{in}$}, with $\mathbf{M}_{up}$} obtained by upsampling $\mathbf{M}_{in}$}. The model output is denoted as $\mathbf{H}_{out}$}, and the corresponding ground truth or basic reference features (if available) are represented as $\mathbf{H}_{gt}$}, the loss of pan-sharpening $\mathbf{L}_{pan}$} can be articulated as follows:
$\mathbf{\mathcal{L}_{pan}}={{\left\|{\mathbf{H}_{gt}}-{\mathbf{H}_{out}}\right\|}_{1}}$ (15)
In the MIF task, we utilize two input images and the fused image to compute a composite loss function $\mathbf{L}_{fus}$} comprising intensity loss $\mathbf{L}_{1}$}, structural similarity index (SSIM) loss $\mathbf{L}_{ssim}$}, and gradient loss $\mathbf{L}_{grad}$}.
${\mathbf{L}_{1}}={{\left\|{\mathbf{H}_{out}}-\max({\mathbf{M}_{up}},{\mathbf{P}_{in}})\right\|}_{1}}$ (16)
${\mathbf{L}_{ssim}}=0.5\times\mathbf{s}({\mathbf{H}_{out}},{{M}_{up}})+0.5\times\mathbf{s}({\mathbf{H}_{out}},{{P}_{in}})$ (17)
$gra{{d}_{join}}=\max(\mathbf{sobel}({\mathbf{M}_{up}}),\mathbf{sobel}({\mathbf{P}_{in}}))$ (18)
${\mathbf{L}_{grad}}={{\left\|\mathbf{sobel}({\mathbf{H}_{out}})-gra{{d}_{join}}\right\|}_{1}}$ (19)
${\mathbf{L}_{fus}}=20{\mathbf{L}_{1}}+50{\mathbf{L}_{ssim}}+100{\mathbf{L}_{grad}}$ (20)
Here, $\mathbf{s}$ denotes the structural similarity index (SSIM) calculation operation on the image, while $\mathbf{sobel}$ represents the Sobel convolution operation.
TABLE II: Evaluation of our method on real-world full-resolution scenes from the WV2 dataset. Bold and underline show the best and second-best values, respectively.
Metric SFIM Brovey GFPCA INNformer SFINet PanFlowNet FAME DISPNet Pan-mamba ARConv Ours
$D_{\lambda}$ $\downarrow$ 0.1403 0.1026 0.1139 0.0995 0.1034 0.0966 0.0951 0.0944 0.0966 0.0963 0.0941
$D_{S}$ $\downarrow$ 0.1320 0.1409 0.1535 0.1305 0.1305 0.1274 0.1263 0.1264 0.1272 0.1276 0.1266
QNR $\uparrow$ 0.7826 0.7728 0.7532 0.7858 0.7827 0.7910 0.7933 0.7938 0.7911 0.7910 0.7939
@@ III-E Loss Function
Figure 6: Comparative visual experiments of several methods on WV2 datasets
Figure 7: Comparative visual experiments of several methods on GF2 datasets
Figure 8: The result of our approach on full-resolution WV2 dataset.
IV Experiments
IV-A Datasets and Benchmark
For the pan-sharpening task, we utilized datasets from WorldView-II (WV2), GaoFen2 (GF2), and WorldView-III (WV3), encompassing a diverse range of urban and natural scenes. Training samples were prepared using the Wald protocol [51], which is widely adopted when the ground truth is inaccessible. To evaluate the effectiveness of our proposed method, we conducted a comprehensive comparison against both traditional and deep learning-based approaches. The baseline methods included classical algorithms such as SFIM [34], Brovey [7], GS [19], IHS [11], GFPCA [28], SFNLR [58] and LRTCFPan [56],as well as state-of-the-art deep learning techniques, including MSDCNN [68], INNformer [74], SFINet [75], MSDDN [14], PanFlowNet [66], FAME [13], DISPNet [52], Pan-Mamba [12], and ARConv [54].
TABLE III: Evaluation of parameter count and GFLOPs.
Methods SRPPNN INNformer MSDDN PanFlowNet FAME DISPNet Panmamba ARConv Ours
Params(M) 1.7114 0.0706 0.3185 0.0873 0.6286 0.4433 0.1827 4.4147 0.2077
GFLOPs(G) 21.1059 1.3079 2.5085 5.7126 10.1943 7.7659 3.0088 14.0833 3.4021
Infer Time(ms) 6.7071 11.9172 7.4857 5.3810 12.7889 11.5221 19.0202 63.3764 24.4713
TABLE IV: Quantitative comparison of MIF task on three datasets. bold and underline show the best and second-best values.
@@ IV-A Datasets and Benchmark
EMFusion 0.943 0.685 0.783 0.611 1.190 0.552 0.475 0.633 0.885 0.665 0.692 0.606
MSRPAN 1.017 0.581 0.799 0.591 1.319 0.436 0.455 0.631 0.960 0.525 0.560 0.577
SwinFusion 1.642 0.703 0.683 0.363 1.537 0.522 0.545 0.290 1.678 0.744 0.720 0.342
Zero 0.950 0.635 0.774 0.581 0.768 0.320 0.582 0.600 1.046 0.582 0.681 0.590
U2Fusion 0.947 0.460 0.292 0.247 0.309 0.074 0.489 0.021 0.865 0.419 0.696 0.240
CDDFuse 1.481 0.650 0.765 0.614 1.589 0.526 0.530 0.612 0.995 0.786 0.719 0.585
PSLPT 0.888 0.548 0.373 0.408 0.675 0.502 0.432 0.405 0.850 0.359 0.325 0.467
Fusionmamba 1.399 0.791 0.736 0.618 1.532 0.538 0.587 0.664 1.368 0.761 0.732 0.619
TC-MoA 1.387 0.540 0.615 0.563 1.584 0.451 0.503 0.414 1.487 0.554 0.594 0.557
PRRGAN 1.405 0.641 0.696 0.607 1.477 0.502 0.497 0.563 1.109 0.682 0.701 0.589
TFS-Diff 1.489 0.665 0.758 0.621 1.591 0.530 0.547 0.658 1.549 0.743 0.684 0.611
Ours 1.491 0.797 0.741 0.628 1.580 0.546 0.592 0.665 1.470 0.820 0.747 0.620
TABLE V: Quantitative comparison of IVIF task on three datasets. bold and underline show the best and second-best values.
MSRS RoadScene M3FD
[…]
SemLA 1.392 0.608 0.290 0.498 1.269 0.564 0.415 0.518 1.495 0.542 0.363 0.473
U2Fusion 1.258 0.512 0.391 0.440 1.605 0.564 0.506 0.546 1.753 0.673 0.578 0.463
SwinFusion 1.647 0.825 0.558 0.504 1.576 0.614 0.450 0.534 1.588 0.746 0.616 0.492
CDDFuse 1.549 0.819 0.548 0.459 1.707 0.610 0.450 0.515 1.673 0.802 0.613 0.460
PSLPT 1.374 0.753 0.553 0.501 1.009 0.134 0.171 0.238 0.638 0.958 0.321 0.483
Fusionmamba 1.635 0.974 0.652 0.511 1.322 0.635 0.543 0.519 1.414 0.747 0.580 0.480
TC-MoA 1.661 0.811 0.565 0.515 1.562 0.577 0.477 0.522 1.556 0.579 0.508 0.466
MLFuse 1.520 0.753 0.519 0.498 1.595 0.629 0.527 0.545 1.600 0.592 0.460 0.501
Up-Fusion 1.584 0.813 0.602 0.502 1.718 0.644 0.536 0.533 1.616 0.816 0.621 0.496
Ours 1.663 0.983 0.655 0.526 1.708 0.651 0.540 0.548 1.766 0.952 0.619 0.515
In the MIF task, we employed medical images from the Harvard Medical website, comprising pairs such as MRI-CT, MRI-PET, and MRI-SPECT [40, 55]. For this task, we compared our method with various deep learning-based multi-modal fusion techniques, including EMFusion [62], MSRPAN [6], SwinFusion [36], Zero [20], U2Fusion [60], CDDFuse [71], PSLPT [53], Fusionmamba [43], TC-MoA [77], PRRGAN [15], and TFS-Diff [65].
@@
Figure 10: Comparative visual experiments of several methods on MRI-SPECT datasets
IV-B Implement Details
All experiments were conducted using the PyTorch framework on two NVIDIA RTX 4090 GPUs. We trained for 500 epochs with a batch size of 4 for the pan-sharpening task and 200 epochs with a batch size of 1 for the medical image fusion (MIF) task. The network parameters were optimized using the Adam optimizer. The initial learning rate was set to 5e45e-4, which was subsequently reduced to 5e85e-8 using the CosineAnnealingLR scheduler over the specified epochs. For the pan-sharpening task, we randomly cropped training set images to obtain LRMS patches of size 32x32 and PAN images of size 128x128. For the MIF task, the training set images were cropped to 256x256.
Figure 11: Comparative visual experiments of several methods on MRI-CT datasets
TABLE VI: Ablation for our methods on three datasets. The PSNR/SSIM/SAM/ERGAS values on benchmarks are reported. The best results are shown in bold. RSS denotes Random-Shuffle Scanning.
@@ IV-B Implement Details
Single Operation (1) RM w/o RSS 42.1460/0.9723/0.0211/0.9042 47.4284/0.9890/0.0102/0.5355 31.3915/0.9329/0.0676/2.8113
(2) RCIM w/o RSS 42.2443/0.9729/0.0210/0.8919 47.6847/0.9899/0.0099/0.5292 31.3476/0.9325/0.0680/2.8230
(3) RMIM w/o RSS 42.2136/0.9727/0.0211/0.8978 47.5787/0.9895/0.0100/0.5324 31.0971/0.9295/0.0704/2.9097
Multi Operations (1) + (2) 42.1205/0.9719/0.0214/0.9075 47.4000/0.9878/0.0104/0.5405 31.2150/0.9316/0.0692/2.8782
(1) + (3) 42.1128/0.9719/0.0215/0.9086 47.3851/0.9876/0.0105/0.5420 31.0547/0.9292/0.0706/2.9053
(2) + (3) 42.1684/0.9721/0.0213/0.9037 47.4655/0.9889/0.0103/0.5381 31.1200/0.9294/0.0702/2.9019
(1) + (2) + (3) 42.0765/0.9718/0.0218/0.9094 47.3111/0.9857/0.0107/0.5498 31.0026/0.9274/0.0719/2.9142
TABLE VII: Ablation of different scanning strategies. The best results are shown in bold. RSS denotes Random-Shuffle Scanning, $S$ denotes Sequential Scanning, BD denotes Bidirectional Scanning, and DS denotes Diagonal Scanning
WorldView-II
[…]
RSS -> SS 42.0967 0.9718 0.0218 0.9094
RSS -> BS 42.2294 0.9726 0.0211 0.8933
RSS -> DS 42.1539 0.9721 0.0214 0.8967
Ours (RSS) 42.3428 0.9734 0.0208 0.8840
IV-C Comparison with SOTA Methods
IV-C1 Pan-sharpening
@@
The experimental results on three datasets are presented in Table I. Reference metrics, including PSNR, SSIM, SAM, and ERGAS, are used to evaluate the fusion effect. The results demonstrate that the proposed method outperforms the SOTA methods across all metrics. Notably, in the PSNR metric, our method achieves improvements of 0.1047, 0.2727, and 0.2301 dB over Pan-mamba, which has the second-best performance. Better PSNR and SSIM evidence that the fusion results transfer more information from the original image and experience less distortion. In the qualitative comparison, Figure 5, Figure 6, and Figure 7 present representative samples from the respective datasets. Our method demonstrates superior performance in the mean squared error (MSE) graph, indicating that the fusion results are closer to the ground truth. Compared to other approaches, our method achieves more accurate restoration of both spectral and spatial details, underscoring the effectiveness of our fusion technique.
@@
To further verify the generalization ability of our method in full-resolution scenes, we use three non-reference metrics, including ${D}_{s}$}, ${D}_{\lambda}$}, and QNR, to evaluate our method on the full WorldView-II dataset. The metric results are summarized in Table II, while Figure 8 provides a visual comparison. Our method consistently outperforms the competing approaches across all metrics, highlighting the robust adaptability of Shuffle Mamba in image fusion tasks.
To comprehensively evaluate the parameter efficiency and computational performance of the proposed method, we further report additional metrics, including the number of parameters, GFLOPs, and actual inference time. All methods were benchmarked on $128\times 128$ resolution images from the WorldView-II dataset, and the results are summarized in Table III. Our method achieves a comparable parameter count to recent state-of-the-art (SOTA) approaches, while being significantly more lightweight—requiring only one-third to one-half the parameters of FAME [13] and DISPNet [52]. Moreover, it consistently outperforms the latest ARConv [54] method in terms of parameter count, computational cost, and inference time. As for training efficiency, due to substantial differences in model structures, we provide a focused comparison between Pan-mamba [12] and our method under the same number of epochs and settings. Pan-mamba completes training in 4.6301 hours (100%), while our method takes 5.2194 hours (113%). This indicates a modest 13% increase in training time. However, our method achieves consistently superior performance across both quantitative and qualitative evaluations, demonstrating its effectiveness in enhancing image fusion quality. Overall, these results highlight the efficiency and practicality of the proposed Shuffle Mamba framework, offering a compelling trade-off between performance and computational cost.
IV-C2 Medical Image Fusion